# StatLLaMA: A multi-stage training framework for building a domain-optimized statistical language model

### v1

@@ 6.2 Limitations
Fourth, the DTFT stage, though informative, remains preliminary. While low-intensity fine-tuning proved essential, we did not exhaustively explore the space of parameter-efficient fine-tuning (PEFT) configurations, hyperparameters, or data compositions. Our investigation into statistical conversation data was also incomplete, leaving open questions regarding optimal integration strategies and effects on real-world conversational performance.
@@ 6.3 Future directions
Advances in fine-tuning and PEFT techniques also hold potential. Comparing LoRA variants, exploring AdapterFusion [Pfeifferetal2021adapterfusion], or partially freezing modules may produce better stability–adaptability trade-offs.
@@ Acknowledgments
Table A2: Full set of prompt templates used in data preparation. Each dataset is generated by Gemini using its corresponding prompt, encompassing diverse statistical reasoning and dialogue tasks to ensure comprehensive understanding across training strategies.
@@
Table A2: Full set of prompt templates used in data preparation. Each dataset is generated by Gemini using its corresponding prompt, encompassing diverse statistical reasoning and dialogue tasks to ensure comprehensive understanding across training strategies.

### v2

@@ 6.2 Limitations
Fourth, the DTFT stage, though informative, remains preliminary. While low-intensity fine-tuning proved essential, we did not exhaustively explore the space of PEFT configurations, hyperparameters, or data compositions. Our investigation into statistical conversation data was also incomplete, leaving open questions regarding optimal integration strategies and effects on real-world conversational performance.
@@ 6.3 Future directions
Advances in fine-tuning and PEFT techniques also hold potential. Comparing LoRA variants, exploring AdapterFusion [48], or partially freezing modules may produce better stability–adaptability trade-offs.
@@ Acknowledgments
Table 12: Full set of prompt templates used in data preparation. Each dataset is generated by Gemini using its corresponding prompt, encompassing diverse statistical reasoning and dialogue tasks to ensure comprehensive understanding across training strategies.
@@
Table 12: Full set of prompt templates used in data preparation. Each dataset is generated by Gemini using its corresponding prompt, encompassing diverse statistical reasoning and dialogue tasks to ensure comprehensive understanding across training strategies.